OpenStudy (anonymous):

Find the sum of the 50 greatest negative integers.

OpenStudy (anonymous):

-1 + -2 + -3 + ....... + -50 =

OpenStudy (anonymous):

I have to use arithmetic series.. and i don't get the question

OpenStudy (anonymous):

-50(25) + -25 = -1275

OpenStudy (anonymous):

which formula are you using?

OpenStudy (anonymous):

Sn = n/2 [2a + (n-1)d]

OpenStudy (anonymous):

n = 50 because there are 50 terms a = -50 because my first term is -50 d = 1 because the difference between -50 and -49 is 1 Therefore, Sn = (50/25)[2(-50 + (50 -1)(1)] = -1275